In a collision of sufficient force, automobile air bags respond by electrically triggering the explosive decomposition of sodium azide (NaN3) to its elements. A 33.4−g sample of sodium azide was decomposed, and the nitrogen gas generated was collected over water at 26°C. The total pressure was 757.0 mmHg. How many liters of dry N2 were generated? Do not answer in scientific notation.

1
Expert's answer
2021-03-10T06:05:22-0500

The decomposition of sodium azide, 2 moles, produce 3 moles of Nitrogen gas. If you do stoichiometry, 33.4 grams of Sodium azide produces 0.771 moles of Nitrogen gas. Use the equation PV= NRT Volume =? n= 0.771 R= 0.082 T =273 + 26 = 299 P = 757 - 25.2 = 731. Eight torrs You need to subtract 25.2 because it is the vapour pressure of water, and you need the volume of dry Nitrogen. The vapour pressure or aqueous tension of water at 26 degrees Celsius is 25.2. Now convert the pressure in atm. Pressure in mm of Hg must be converted into atm by dividing with 760 then the pressure in atm would be 0.963 0.963 atm * V = 0.771 moles * 0.082 lit- atm/Kel-mol * 299 Solving for Volume in litres should be 19.63 litres